OpenStudy (jacobtaylorr99):

Determine the solution set of (3x + 1)2 - 100 = 0.

OpenStudy (mathmale):

Hello, Jacob!, Suggest writing (3x + 1)2 - 100 = 0. as (3x + 1)^2 - 100 = 0. The ^ symbol calls for exponentiation. 1. Add 100 to both sides, to move the 100 to the right side.

OpenStudy (mathmale):

(3x + 1)^2 = 100

OpenStudy (mathmale):

2.

OpenStudy (mathmale):

Take the square root of both sides, remembering to put a "plus or minus" sign in front of

OpenStudy (mathmale):

he

OpenStudy (mathmale):

the square root of 100. What do you have now? This equation may have two solutions. It's important that you check them (by substitution of your result into the original equation).

OpenStudy (triciaal):

Welcome to Openstudy111 you can shift the number 6 key to get the symbol for exponents follow the steps given by @mathmate

OpenStudy (mathmate):

@triciaal haha, I suppose you mean @mathmale, my virtual twin, lol.